A sample of size 39 will be drawn from a population with mean 23 and standard deviation 10. Find the probability that  will be greater than 25.


1
Expert's answer
2021-03-16T07:41:27-0400

Let "X" be a random variable of a sample of size 39 from a population which is normally distributed with mean "23" and standard deviation "10."

Then "\mu =23" and "\sigma =10"

Let "Z=\frac {X-\mu}{\sigma}." Then "Z=\frac{X-23}{10}".

We have to find "P(X>25)."

"\therefore P(X>25)=P(Z>\frac{25-23}{10})"

"=P(Z>0.2)"

"=0.5-P(0\leq Z\leq0.2)"

"=(0.5-0.0792)" "=0.42"

So the required probability is "=0.42"